# A Simple Unified Uncertainty-Guided Framework for Offline-to-Online Reinforcement Learning

### v2

@@ D.1 Experiments on AntMaze
Table 6: Comparison of the averaged D4RL score on AntMaze tasks with CQL as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
CQL offline-ft online-ft BR O3F APL PEX SUNG
antmaze-umaze-v2 88.6 87.9
$\pm$1.0
96.0
$\pm$1.1
68.5
$\pm$3.5
98.6 $\pm$0.5 96.0
$\pm$1.6
87.3
$\pm$2.0
96.9
$\pm$0.6
antmaze-umaze-diverse-v2 41.0 45.9
$\pm$1.8
0.0
$\pm$0.1
0.0
$\pm$0.0
0.0
$\pm$0.0
0.0
$\pm$0.0
0.7
$\pm$1.4
50.5 $\pm$3.1
antmaze-medium-play-v2 63.8 75.4
@@
18.6
$\pm$11.0
22.2
$\pm$8.2
67.3
$\pm$2.3
22.8
$\pm$2.5
0.3
$\pm$0.4
86.3 $\pm$1.4
antmaze-medium-diverse-v2 61.8 74.1
$\pm$1.8
22.1
$\pm$9.1
13.6
$\pm$4.9
80.8
$\pm$2.9
36.8
$\pm$4.3
0.3
$\pm$1.0
85.6 $\pm$2.1
antmaze-large-play-v2 32.0 41.8
$\pm$3.7
0.1
$\pm$0.2
0.0
$\pm$0.0
0.0
$\pm$0.0
0.0
$\pm$0.0
0.0
$\pm$0.0
52.7 $\pm$2.8
antmaze-large-diverse-v2 32.6 35.3
$\pm$2.6
0.7
$\pm$1.8
0.1
$\pm$0.1
0.0
$\pm$0.0
0.0
$\pm$0.0
0.0
$\pm$0.0
44.1 $\pm$3.5
Total 319.6 360.4 137.5 104.3 246.7 155.6 88.6 416.1
Table 7: Comparison of the averaged D4RL score on MuJoCo tasks, with a focus on both final performance and online improvement $\delta$. m = medium, m-r = medium-replay. We run 200K environment steps to keep consistent with ODT’s setting. We report the mean value over 5 seeds.
ODT (offline) ODT (online) $\delta$ TD3+BC SUNG $\delta$ CQL SUNG $\delta$
halfcheetah-m-v2 42.3 42.1 -0.2 48.0 85.5 37.5 48.1 86.7 38.6
hopper-m-v2 62.3 89.8 27.5 61.5 97.1 35.6 73.7 106.1 32.4
walker2d-m-v2 72.1 73.7 1.6 82.2 115.4 33.2 84.3 114.6 30.3
halfcheetah-m-r-v2 37.7 40.6 2.9 44.6 79.8 35.2 46.9 82.9 36.0
hopper-m-r-v2 83.2 88.3 5.1 55.9 104.1 48.2 96.0 83.5 -12.5
walker2d-m-r-v2 67.8 73.1 5.3 71.7 110.7 39.0 83.4 116.7 33.3
Total 365.4 407.6 42.2 364.9 592.6 228.7 432.4 590.5 158.1
We conduct experiments on more challenging AntMaze domains, which require the agents to deal with the sparse rewards and stitch fragments of sub-optimal trajectories. Results for AntMaze domains with SPOT and CQL as backbone offline RL method are shown in Table 5 and 6, respectively. From the results, we can observe that online-ft, BR and PEX struggle to benefit from online finetuning in most settings, except for the relatively easy antmaze-umaze task. This stems from the need for explicit or implicit behavior cloning that is offered by offline RL objectives, which is required to handle the sparse rewards and stitch sub-optimal sub-trajectories. However, online-ft, BR and PEX are developed on top of purely online RL methods, thus failing in most of the tasks.
Furthermore, we remark that SUNG, when combined with either SPOT or CQL, outperforms other state-of-the-art methods in most settings. This demonstrates the superiority of SUNG for discovering informative state-action pairs that can lead to high outcomes, thereby providing valuable behavior patterns for agents to learn from. Moreover, SUNG achieves flexible trade-off between offline and online RL objectives by tuning $p$, which enables stitching sub-optimal sub-trajectories and overcoming the sparse reward issue.
D.2 Comparison with Online Decision Transformer
Since online decision transformer (ODT) Zheng et al. [2022] is built on top of upside-down RL Schmidhuber [2019], Chen et al. [2021], Emmons et al. [2022], we compare SUNG with ODT in this separate subsection. We follow the official implementation in https://github.com/facebookresearch/online-dt for reproduction. We strictly follow the original experimental setting in ODT to take 200K environment steps and only consider tasks with medium and medium-replay offline datasets for online finetuning to achieve fair comparison.
As shown in Table 7, for almost all the settings, SUNG when combined with either TD3+BC or CQL outperforms ODT by a large margin in both final performance and online improvement $\delta$. This is because ODT learns a reward-conditioned policy in a supervised learning paradigm, thereby struggling to fully benefit from the trail-and-error paradigm used in traditional online RL. However, it is worth noting that ODT enables users to customize the behavior of the agents by specifying different conditions, which is not featured by traditional RL.
D.3 Learning Curves
@@
D.4 Ablation Study
Ablation Study with Variants of SUNG (Fig. 2). We consider the following variants of SUNG to perform the ablation study over the components of our framework:
  • w/o Opt. Exploration: We replace the proposed optimistic exploration strategy by the default exploration provided by TD3 and SAC. For TD3, we select action with Gaussian noise for exploration, i.e., $\pi_{E}(s)=\pi_{\phi}(s)+\epsilon,\epsilon\sim\mathcal{N}(0,\delta)$. For SAC, we sample action from the policy for exploration, i.e., $\pi_{E}(s)=a,a\sim\pi_{\phi}(\cdot|s)$.
  • Opt. Exploration w/o Unc.: We remove the uncertainty from the proposed exploration strategy to greedily select action that maximizes Q value, i.e., $\pi_{E}(s)=\arg\max_{a\in\mathcal{C}}Q(s,a)$.
  • Opt. Exploration w/o Q: We remove the Q value from the proposed exploration strategy to greedily select action that maximizes the uncertainty, i.e., $\pi_{E}(s)=\arg\max_{a\in\mathcal{C}}\mathcal{U}(s,a)$.
  • w/o Adp. Exploitation: We replace the proposed adaptive exploitation method by the standard online RL objective of TD3 and SAC, respectively.
  • Unc. Estimation w/ Q std: We replace VAE by the standard deviation of two Q value functions for uncertainty estimation, i.e.,
    $\mathcal{U}(s,a)=\sigma(Q_{\theta_{1}}(s,a),Q_{\theta_{2}}(s,a))=|Q_{\theta_{1}}(s,a)-Q_{\theta_{2}}(s,a)|.$
Besides, we also consider two more ablation settings:
@@ D.4 Ablation Study
Hyper-parameter Analysis on Softmax Temperature $\alpha$. Here, we give a hyper-parameter analysis of the softmax temperature $\alpha$ in Eq. (8). Specifically, we investigate the choice of $\alpha\in\{0.1,0.2,0.5,1.0,5.0,10.0,20.0,100.0\}$ for MuJoCo domains in Fig. 9. As shown in the figure, we can find that $\alpha$ is a relatively insensitive hyper-parameter for the offline-to-online performance, and that the best-performing choice of $\alpha$ varies among different settings. Thus, though a careful tuning on $\alpha$ will bring more offline-to-online improvement, we set $\alpha=1.0$ in all the experimental result throughout the paper to guarantee the simplicity of the proposed framework.
@@
Hyper-parameter Analysis on OORB sampling probability $p_{\mathrm{OORB}}$. Furthermore, we give a hyper-parameter analysis on $p_{\mathrm{OORB}}$ for the simple-yet-effective OORB. Specifically, we investigate the choice of OORB sampling probability $p_{\mathrm{OORB}}\in\{0.0,0.1,0.2,0.4,0.6,0.8,1.0\}$ for MuJoCo domains in Fig. 10. Note that a higher probability $p_{\mathrm{OORB}}$ indicates more reuse of the offline dataset during policy learning. From the results, we can observe that $p_{\mathrm{OORB}}=0.1$ performs the best when combined with either TD3+BC and CQL, while either setting $p_{\mathrm{OORB}}$ too conservatively (i.e., $p_{\mathrm{OORB}}=0$) or too adversely (i.e., $p_{\mathrm{OORB}}=1.0$) hurts the performance. As pointed out by Zheng et al. Zheng et al. [2023], offline data can prevent agents from prematurely converging to sub-optimal policies due to the potential data diversity, while online data can stabilize training and accelerate convergence. Thus, it is crucial for sample-efficient offline-to-online RL algorithms to incorporate both offline and online data during policy learning.

### v3

@@ V-F SUNG with Other Advanced Techniques (RQ4)
TABLE V: Comparison of the averaged D4RL score on MuJoCo tasks with CQL-10 as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
CQL-10 BRPQ
RLPD
(UTD=20)
SUNG
(UTD=1)
SUNG
(UTD=5)
halfcheetah-r-v2 29.9 85.8
$\pm$17.2
72.5
$\pm$5.9
92.7 $\pm$4.2 97.5 $\pm$5.0
hopper-r-v2 7.4 28.4
$\pm$12.0
87.8
$\pm$14.0
62.5
$\pm$13.8
89.7 $\pm$28.4
walker2d-r-v2 21.6 16.1
$\pm$4.0
65.7
$\pm$16.4
33.0
$\pm$19.6
73.1 $\pm$33.0
halfcheetah-m-v2 55.0 80.3
$\pm$20.5
84.2
$\pm$2.2
96.6 $\pm$1.0 105.5 $\pm$3.6
hopper-m-v2 66.9 79.5
$\pm$19.0
98.1
$\pm$12.2
111.4 $\pm$0.8 86.3
$\pm$24.2
walker2d-m-v2 83.2 74.4
$\pm$13.9
114.3 $\pm$2.2 113.8
$\pm$2.6
111.3
$\pm$25.3
halfcheetah-m-r-v2 52.6 74.8
$\pm$20.7
80.3
@@
92.5 $\pm$0.5 96.7 $\pm$6.1
hopper-m-r-v2 102.3 71.6
$\pm$10.1
75.0
$\pm$17.9
100.7 $\pm$13.6 64.9
$\pm$14.8
walker2d-m-r-v2 82.1 71.7
$\pm$11.2
108.5
$\pm$2.8
105.1
$\pm$14.6
117.3 $\pm$5.8
Total 501.0 582.6 786.4 808.2 842.3
Avg Improvement / 16.28% 56.96% 61.31% 68.12%
In this subsection, we aim to illustrate that SUNG can be seamlessly compatible with other advanced RL techniques, leading to further improved finetuning performance. Particularly, we focus on: (1) Ensemble Q learning, as demonstrated in [55, 6], exhibits robust performance in offline pretraining and online finetuning. Nevertheless, it is crucial to acknowledge that the ensemble mechanism introduces additional computational costs. (2) High update-to-data (UTD) ratio [56, 22], as discussed in [56, 22], enhances sample efficiency by maximizing the number of gradient updates per environment step. Nonetheless, this efficiency comes at the expense of increased computational costs.
Specifically, we follow previous works [6, 22] to utilize CQL-10, a variant of CQL employing ensemble Q learning with 10 Q functions, for offline pretraining. We compare SUNG against two baselines: (1) BRPQ [6] introduces both prioritized replay buffer and pessimistic Q ensembles for offline-to-online RL. (2) RLPD [22] is a recent state-of-the-art method that utilizes offline data to accelerate online RL, benefiting from both the ensemble technique and high UTD ratio. For a fair comparison, we present results of SUNG with UTD ratio of both 1 and 5.
We present the experimental results in Table V. We can observe that SUNG employing the ensemble Q learning technique, indicated by UTD=1, exhibits a significant performance advantage over BRPQ, showcasing an additional 45.03% improvement in offline-to-online performance. In contrast, SUNG, when combined with both ensemble Q learning and a high UTD ratio, denoted as UTD=5, attains results that are either superior or, at the very least, comparable to those of RLPD, yielding an additional 11.16% improvement in offline-to-online performance. The findings above underscore the seamless synergy of SUNG with other advanced RL techniques for enhanced finetuning performance.
Furthermore, we remark that SUNG with UTD ratio of 1 exhibits weak results for hopper-r and walker2d-r, suggesting that poorly initialized policies are challenging to recover through low UTD ratio online finetuning. In contrast, SUNG with UTD ratio of 5 demonstrates competitive results in these settings. However, we notice that SUNG with UTD ratio of 5 shows subpar performance in hopper-m and hopper-m-r, possibly due to the statistical overfitting [56]. As suggested in previous works, regularization approaches, such as simple L2 normalization, layer normalization, dropout and random ensemble distillation, can eliminate this issue. We consider this avenue for future work.
TABLE VI: Comparison of training time with CQL as the offline RL backbone method.
offline-ft online-ft BR APL PEX PROTO
SUNG
w/ VAE
SUNG
w/ Q-Ensemble
1h 1h 0.8h 1h 0.6h 0.8h 0.7h 2.4h
V-G Comparison on Training Time
In this subsection, we compare the training time of various offline-to-online RL methods. The results are summarized in Table VI, where we use CQL as the backbone method and report the time required for 100K environment steps in the halfcheetah-m-v2 setting. The experiments are conducted on a single NVIDIA RTX 3090 GPU and Intel Xeon Platinum 8255C CPU at 2.50GHz. Notably, we include SUNG w/ Q-Ensemble, which utilizes 10 Q-networks for uncertainty estimation. As shown in the results, all baseline methods require approximately 1 hour of training, except for SUNG w/ Q-Ensemble, which takes 2.4 hours. This also highlights the computational efficiency of our proposed VAE-based uncertainty estimation methods.
VI Conclusion
This paper investigates the effective finetuning methods of pretrained offline RL agents via online interactions. We introduce SUNG, a simple yet effective uncertainty-guided framework, designed to address the challenges of constrained exploratory behavior and state-action distribution shift. SUNG employs a VAE estimator to quantify state-action uncertainty. The optimistic exploration strategy, facilitated by a bi-level action selection mechanism, enables SUNG to efficiently explore the state-action space. Furthermore, SUNG incorporates an adaptive exploitation method with OOD sample identification, creating a seamless bridge between the offline and online learning stages. Extensive empirical results, spanning various backbone offline RL methods, environments, and datasets, confirm the superiority of SUNG. Notably, SUNG demonstrates robust finetuning performance across diverse hyperparameter settings and can be seamlessly integrated with different advanced RL techniques to achieve further improved finetuning results.
VII Limitation and Future Work
While achieving promising performance, the proposed framework also has some inherent limitations. SUNG investigates the offline-to-online RL setting to solve tasks whose inputs are the states of the simulated environment. It would be interesting to see whether SUNG is also superior with the visual RL setting. We also remark that the exploration of its application on real-world tasks is also an interesting future work. Furthermore, SUNG may struggle in finetuning pretrained agents from offline datasets with low-quality behaviors, highlighting the need for more robust policy improvement methods within offline-to-online RL algorithms. Finally, further investigation on uncertainty estimation, replay buffer, and other backbone offline RL methods needs to be explored.
Appendix A Further Discussion on Other Uncertainty Quantification Methods
In this section, we provide more detailed discussion with other uncertainty quantification methods in RL, where we mainly categorize them as (1) Distributional RL, (2) Ensemble based methods, (3) Bootstrapping based methods, (4) Monte Carlo dropout based methods, and (5) Generative models based methods.
  • Distributional RL, such as C51 [41], QR-DQN [42], and IQN [43], aims to model the aleatoric uncertainty on return distribution that originates from inherent stochasticity of the environment and agent interactions in RL. Different from them, we aim to model epistemic uncertainty that originates from the insufficient knowledge of the environment to guide exploration and exploitation in offline-to-online RL. Thus, the investigated problems of them are inherently different, and the uncertainty quantification methods in distributional RL cannot be applied to general offline-to-online RL setting.
  • Ensemble based methods, such as SUNRISE [13], utilize the variance of the predicted Q values from the ensemble Q networks as the uncertainty. However, their applicability to the offline-to-online RL setting is limited, as they require the offline pretraining phase to also employ ensemble-based offline RL algorithms. In contrast, the VAE-based uncertainty quantification method in SUNG can be generally applicable. We also provide an empirical comparison in Appendix E.A to highlight the effectiveness of the VAE-based method.
  • Bootstrapping based methods, such as Bootstrapped DQN [44], estimate the diversity of multiple heads of Q networks that are trained from different sample subsets as uncertainty. Similar to ensemble based methods, they also require the offline pretraining phase to employ bootstrapping-based offline RL algorithms, which limit their applicability in the offline-to-online RL setting. We do not include the empirical comparison of this method due to its similarity to ensemble-based methods.
  • Monte Carlo dropout based methods, such as UWAC [16], quantify epistemic uncertainty by the variance of Q values from multiple forward passes of the dropout-enabled Q network. Since modern offline RL algorithms typically do not utilize the dropout technique, the applicability of this method is also limited in the offline-to-online RL setting. We try to make comparison between the proposed VAE method and Monte Carlo dropout based methods; however, as modern offline RL algorithms typically utilize shallow neural networks (e.g., three-layer MLP in TD3+BC and CQL), the dropout technique makes the training process quite unstable in practice. We leave the exploration of stable Monte Carlo dropout based offline-to-online RL methods as future work.
  • Generative model based methods, such as CPQ [46], estimate the action density of the behavior policy distribution of the offline dataset as the uncertainty to penalize offline RL polices in generating OOD actions. In contrast, we utilize VAE to estimate density of state-action pairs as uncertainty quantification, which can guide both exploration and exploitation in offline-to-online RL. Although diffusion models [47, 48] are regarded as stronger generative models than VAEs, they require higher computational costs for training and inference. Moreover, the hyper-parameter tuning for diffusion models is non-trivial. As such, we adopt VAE in this work and leave the extension to diffusion models as future work.
Figure 7: Hyper-parameter analysis on $L$ of SUNG combined with TD3+BC and CQL, where a larger $L$ means lower bias on uncertainty quantification.
Appendix B Discussion on bias in Density Estimation
As stated earlier, there theoretically exists a bias between ELBO and the negative log likelihood of the state-action density, since we have $\log p_{\psi}(s,a|z)=\mathcal{L}_{\mathrm{ELBO}}-D_{\mathrm{KL}}\left[q_{\varphi}(z|s,a)||p(z)\right]$. To achieve further bias reduction, we follow [57, 28] to utilize the importance sampling technique to derive the density estimation as:
$\log p(s,a)$ $=\mathbb{E}_{q_{\varphi}(z|s,a)}\left[\log\frac{p_{\psi}(s,a,z)}{q_{\varphi}(z|s,a)}\right]$ (13)
$\approx\mathbb{E}_{q_{\varphi}(z^{(l)}|s,a)}\left[\log\frac{1}{L}\sum_{l=1}^{L}\frac{p_{\psi}(s,a,z^{(l)})}{q_{\varphi}(z^{(l)}|s,a)}\right].$
As theoretically shown in [57], this estimator is the lower bound of the log likelihood of the state-action density; the bound becomes tighter as $L$ increases and it converges to exact equality in the limit as $L\rightarrow\infty$. Particularly, when $L=1$, it is exactly the ELBO estimator.
Now, we give empirical analyses on the hyper-parameter $L$ to demonstrate how the bias in uncertainty quantification impacts the performance of SUNG. We present results of $L\in[1,5,10,20,100]$ in Fig. 7, where a larger $L$ means lower bias on uncertainty quantification. We can find that all the variants yield similar performance, thus the slight bias in uncertainty quantification would not significantly impact the performance of SUNG. As such, we set $L=1$ (i.e., ELBO estimator) in this paper, due to its simplicity and effectiveness.
Figure 8: Comparison of VAE usage in offline RL and offline-to-online RL.
Appendix C More Implementation Details of VAE
We present more implementation details for the VAE here. As the VAE is trained to estimate the state-action density in our work, we design its architecture of as an encoder of a three-layer MLP ($d_{\mathcal{S}}+d_{\mathcal{A}}$, 750, 750, $2(d_{\mathcal{S}}+d_{\mathcal{A}})$) and a decoder of a three-layer MLP ($2(d_{\mathcal{S}}+d_{\mathcal{A}})$, 750, 750, $d_{\mathcal{S}}+d_{\mathcal{A}}$), where $d_{\mathcal{S}}$ and $d_{\mathcal{A}}$ denote the state dimension and action dimension. We illustrate the VAE architecture in Fig. 9. During the training of VAE, the encoder encodes the state-action pair $(s,a)$ and outputs the mean vector $\mu$ and standard deviation $\sigma$, which specify the normal distribution $\mathcal{N}(\mu,\sigma^{2})$ for the hidden representation $z$. Since the sampling process is non-differentiable, we apply the reparameterization trick to sample $z\sim\mathcal{N}(\mu,\sigma^{2})$ by sampling a standard normal distribution $\epsilon\sim\mathcal{N}(0,I)$ and mapping to $\mathcal{N}(\mu,\sigma^{2})$ as $z=\mu+\sigma\odot\epsilon$. Then, the hidden representation is decoded by the encoder as the generated state $\hat{s}$ and action $\hat{a}$. As such, the parameters of the VAE can be optimized by maximizing the ELBO. During the inference of VAE, we only perform the forward pass to derive the ELBO as the uncertainty estimation.
Figure 9: Illustration of the VAE architecture.
Appendix D Full Learning Curves
@@
Appendix E Additional Empirical Analyses
E-A Analyses on Other Uncertainty Quantification Method
In this subsection, we examine an alternative uncertainty quantification method, Bayesian ensembling [58, 59], which estimates epistemic uncertainty by computing the predictive variance across an ensemble of neural networks, each representing a posterior sample obtained via Markov Chain Monte Carlo (MCMC). Specifically, we train CQL [10] with 10 Q networks for offline pretraining, denoted as CQL-10. Then, we consider two variants of SUNG: (1) SUNG w/ VAE and (2) SUNG w/ BE. These variants differ in their uncertainty quantification methods. The former employs a VAE for state-action density estimation, while the latter uses the variance of 10 Q functions to estimate epistemic uncertainty. As demonstrated in Table VII, the proposed VAE-based uncertainty quantification method can attain comparable performance to using the variance of 10 Q functions, which would significantly increase computational costs. Furthermore, the results also demonstrate the potential of the proposed SUNG built on top of other uncertainty quantification methods.
TABLE VII: Comparison of the averaged D4RL score on MuJoCo tasks with CQL-10 as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
CQL-10
SUNG
w/ VAE
SUNG
w/ BE
halfcheetah-m-v2 55.0 96.6
$\pm$1.0
98.8 $\pm$1.3
hopper-m-v2 66.9 111.4
$\pm$0.8
112.5 $\pm$1.1
walker2d-m-v2 83.2 113.8 $\pm$2.6 110.2
$\pm$9.5
halfcheetah-m-r-v2 52.6 92.5
$\pm$0.5
96.6 $\pm$1.0
hopper-m-r-v2 102.3 100.7
$\pm$13.6
107.4 $\pm$4.8
walker2d-m-r-v2 82.1 105.1
$\pm$14.6
113.3 $\pm$16.8
Total 501.0 620.0 638.8
E-B Additional Ablation Studies
We consider two more ablation settings:
@@ E-B Additional Ablation Studies
E-C Hyper-parameter Analysis on Softmax Temperature $\alpha$
Here, we give a hyper-parameter analysis of the softmax temperature $\alpha$ in Eq. (10). Specifically, we investigate the choice of $\alpha\in\{0.0,0.1,0.2,0.5,1.0,5.0,10.0\}$ for MuJoCo domains in Fig. 14. Note that $\alpha=0$ can be regarded as an ablation variant with the greedy selection strategy. As shown in the figure, we can find that $\alpha$ is a relatively insensitive hyper-parameter for the offline-to-online performance, and that the best-performing choice of $\alpha$ varies among different settings. Thus, though a careful tuning on $\alpha$ will bring more offline-to-online improvement, we set $\alpha=1.0$ in all the experimental result throughout the paper to guarantee the simplicity of the proposed framework. Furthermore, when we set $\alpha=0$, the performance drastically decreases, highlighting the necessity of the randomness for sampling in the exploration policy.
@@ E-C Hyper-parameter Analysis on Softmax Temperature $\alpha$
E-D Hyper-parameter Analysis on OORB Sampling Probability $p_{\mathrm{OORB}}$
Furthermore, we give a hyper-parameter analysis on $p_{\mathrm{OORB}}$ for the simple-yet-effective OORB. Specifically, we investigate the choice of OORB sampling probability $p_{\mathrm{OORB}}\in\{0.0,0.1,0.2,0.4,0.6,0.8,1.0\}$ for MuJoCo domains in Fig. 15. Note that a higher probability $p_{\mathrm{OORB}}$ indicates more reuse of the offline dataset during policy learning. From the results, we can observe that $p_{\mathrm{OORB}}=0.1$ performs the best when combined with either TD3+BC and CQL, while either setting $p_{\mathrm{OORB}}$ too conservatively (i.e., $p_{\mathrm{OORB}}=0$) or too adversely (i.e., $p_{\mathrm{OORB}}=1.0$) hurts the performance. Offline data can prevent agents from prematurely converging to sub-optimal policies due to the potential data diversity, while online data can stabilize training and accelerate convergence. Thus, it is crucial for sample-efficient offline-to-online RL algorithms to incorporate both offline and online data during policy learning.